# Controllable Generation with Text-to-Image Diffusion Models: A Survey

### v1

@@ 4.1.1 Subject-Driven Generation
As the basic input for text-to-image diffusion models, text plays a crucial role in adapting these models to specific user needs. Textual Inversion (TI)[41] adopts an innovative approach by embedding user-provided concepts into new ’words’ within the text embedding space. This method expands the tokenizer’s dictionary and optimizes additional tokens using a denoising process on provided images. DreamBooth[42] follows a similar path but utilizes low-frequency words (i.e., sks) to represent concepts and additionally updates the parameters of the UNet with a class-specific prior preservation loss to enhance the diversity of generated outputs. The straightforward and adaptable frameworks of TI and DreamBooth establish them as foundational models for numerous subsequent tuning-based methods. Furthermore, Custom Diffusion [45] analyzes weight deviations during the fine-tuning process and discovers the pivotal role of cross-attention layer parameters, particularly key and value projections (i.e., $W^{k}$ and $W^{v}$). This insight leads to a focused update on these projections and the incorporation of extra text tokens and regularization loss for fine-tuning.
Some approaches have been taken to expand the text embedding space, particularly by considering the distinction of each UNet layer[50, 61]. They apply distinct text embeddings across various layers. In contrast, CatVersion [75] diverges from the focus on text embeddings and the UNet’s parameters and advocates for tuning concatenated embeddings within the feature-dense space of the text encoder. Such a method is suggested to be more effective in learning the nuances between a personalized concept and its base class, contributing to the preservation of prior knowledge within the model.
In addition, parameter-efficient tuning (PEFT)[191, 192, 193, 194] plays a pivotal role in personalization methods [63]. Low-rank Adaptation (LoRA)[192] has seen widespread integration across a range of personalization techniques [42, 59, 181, 177, 104]. Besides, Xiang et al. propose ANOVA [52], which opts for the adapter [191], and reveals that placing adapters subsequent to the cross-attention block enhances performance significantly. To facilitate the comprehensive application and evaluation of PEFT in the fine-tuning of diffusion models, LyCORIS [68] has developed an open-source library [^1] https://github.com/KohakuBlueleaf/LyCORIS. This library encompasses a broad spectrum of PEFT methods, including but not limited to LoRA [192], LoHa, and DyLoRA [193]. LyCORIS further introduces a detailed framework for the systematic analysis and assessment of these PEFT techniques, significantly advancing the field of diffusion model personalization.
Moreover, a critical challenge in the realm of personalization is the disentanglement of specific concepts from the provided samples. Numerous studies [62, 59, 81, 65, 74] have identified a common issue where extraneous information becomes intertwined with the intended concept during the customization process, such as inadvertently learning the context surrounding images in subject-driven generation. To effectively isolate and extract the essential concept information from samples, several works [62, 70, 76] have investigated the use of explicit masks. In a similar vein, Disenbooth [59] and DETEX [81] focus on mitigating the influence of background elements in the personalization process. DETEX goes a step further by also aiming to decouple the pose information of subjects from the overall concept. Meanwhile, PACGen [65] employs aggressive data augmentation techniques, transforming the size and location of the personalized concepts within the samples, thereby aiding in the separation of spatial information from the core concept itself.
Also, training diffusion models on small-scale datasets often meet another significant challenge: it risks compromising the generative model’s broader applicability, leading to the requirement of a delicate balance between fidelity and editability [82]. To tackle this issue, several studies have introduced preservation mechanisms, focusing on strategies to prevent overfitting to input samples [42, 45, 58, 64, 51, 78]. For instance, Perfusion [58] addresses this by locking concepts’ cross-attention keys to their prior categories and employing a gated rank-1 method for concept learning. SVDiff [51] takes a different approach, adjusting the singular values in the weight matrices of the model. This technique is designed to minimize overfitting risks and mitigate issues like language drifting. Furthermore, OFT [64] emphasizes the importance of the hyperspherical energy in weight matrices for sustaining the model’s semantic generative capabilities. Hence, it introduces an orthogonal fine-tuning method, further contributing to the preservation of the model’s generalization ability in the face of limited training data.
In addition to the aforementioned methods, several researchers have explored alternative training techniques aimed at optimizing generative performance, expediting the tuning process, and minimizing GPU memory usage [46, 57, 73, 56, 71, 75]. Specifically, DVAR [46] identifies limitations in standard training metrics for assessing the convergence in concept learning and utilize a simple variance-based early stopping criterion, enhancing the efficiency of the fine-tuning process. Gradient-Free Textual Inversion [57] adopts an innovative approach by dividing the optimization process into two parts: dimension reduction in the search space and non-convex, gradient-free optimization in a subspace. This method achieves a significant speed-up in optimization with minimal impact on performance. MATTE [73] delves into the roles of timesteps and UNet’s layers in personalizing various concept categories like color, object, layout, and style, aiming to enhance performance across different concept types. Addressing the need for high-quality data in Textual Inversion, COTI [56] introduces an active and controllable data selection framework to improve Textual Inversion by broadening the data scope. Similarly, He et al. [71] adopt a data-centric approach, proposing a novel strategy for generating regularization datasets at both text and image levels, further enriching the research landscape in this domain.
• Model-based Personalized Score Prediction. Model-based methods employ encoders to embed concepts, offering a significant speed advantage over tuning-based approaches when extracting concepts from images. Some works focus on domain-aware encoders specifically designed to embed images from targeted domains [47, 55]. For instance, InstantBooth [55] employs a specialized encoder and adapters trained on the face and cat domains to extract text embeddings and detailed patch features for concept learning. In contrast, other model-based methods opt for a domain-agnostic approach, training encoders on open-world images to extract more generalized conditions [48, 49, 53, 66, 67, 77, 79, 69]. These methods typically utilize large pre-trained models like CLIP [39] and BLIP-2 [195] as image encoders, focusing on fine-tuning a limited number of parameters, such as a projection layer [48, 49, 82, 66]. ELITE [48], for example, integrates a global mapping network and a local mapping network based on CLIP [39]. The global network transforms hierarchical image features into multiple text embeddings, while the local network infuses patch features into cross-attention layers for detailed reconstruction. BLIP-Diffusion [60] advances customization by pre-training a BLIP-2 [195] encoder for text-aligned image representation and developing a task for learning subject representations, enabling the generation of novel subject renditions. Following on E4T [47], Arar et al. [66] introduce an encoder for acquiring text embeddings and propose a hypernetwork to predict LoRA-style attention weight offsets in UNet. SuTI [53] takes a unique approach inspired by apprenticeship learning [196], training a vast array of expert models on millions of internet image clusters. The apprentice model is then taught to imitate these experts’ behaviors. CAFE [80] build a customization assistant based on pre-trained large language model and diffusion model.
• Training-free Personalized Score Prediction. The pivotal technique for training-free personalization is extracting concept information from reference images in synthesis process. Similar to retrieval-augmented generation in natural language processing, leveraging knowledge from samples helps models faithfully generate given concepts. Re-Imagen [43] represents a novel approach to generating images of uncommon or rare categories, such as Chortai(dog)’ and Picarones(food)’. This method leverages an external multimodal knowledge base, utilizing relevant image-text pairs retrieved from this database as references for image generation. Beyond this, several methods, whether tuning-based or model-based, incorporate the use of reference images to enhance the accuracy and fidelity of the visual details in the generated images [78, 77, 91].
4.1.2 Person-Driven Generation
The person-driven generation task (also known as human-centric personalization) is specifically focused on creating human-centric visual outputs that maintain the same identity as the individuals depicted in the exemplar samples. While person-driven generation is a specialized subset of the broader subject-driven generation category and several methods pertinent to this task have already been discussed in the previous section, we will concentrate on highlighting and analyzing those techniques that are explicitly tailored for person-driven generation in this part.
Similar to model-based subject-driven generation, many person-driven approaches encode facial images into text embedding space to provide identity condition [83, 85, 86, 27, 87]. For example, to achieve a balance between identity preservation and editability, Xiao et al. [83] introduces a novel approach that combines text prompts with visual features derived from reference images of individuals, called FastComposer. Specifically, this method fuses the human-related text embeddings (e.g., ’man’ and ’woman’) with visual features by multilayer perceptron, effectively encapsulating both the textual and visual conditions of the person’s identity. Besides CLIP [39], Face0 [85] and DreamIdentity [86] employs pre-trained face recognition models [197] as their facial image encoders, where Face0 utilizes the Inception ResNet V1 [198] and DreamIdentity introduces a ViT-style [199] Multi-word Multi-scale (M²ID) encoder. While most methods utilize multi-modal pre-trained image encoders (e.g., CLIP [39]) or facial recognition models [197], $\mathscr{W}+$ Adapter [90] introduces an innovative approach using StyleGAN’s inversion encoder.
Inspired by retrieval-augmented generation (RAG), Tang et al. [91] introduce a novel retrieval-based method specifically tailored for human-centric personalization. Complementing this approach, they also present an anime figures dataset, named RetriBooru-V1, which is uniquely characterized by enhanced identity and clothing labels. Central to their method is the use of a frozen Variational Autoencoder (VAE) [200] for encoding reference images and seamless integration into the generation process by cross-attention and zero-convolution layers. These layers play a crucial role in accurately positioning the reference attributes—such as identity and clothing features—at the correct geometric locations in the generated image, thereby ensuring a high degree of fidelity and relevance in the output.
In contrast to subject-driven methods, person-driven generation approaches can benefit significantly from face segmentation, obtained either through parsing models or annotations [83, 27, 95, 96, 97]. For instance, some works, such as Stellar [96], employ face masks to eliminate background elements during data processing, thereby sharpening the focus on human identity within the input data. Conversely, other approaches leverage face masks for constructing [83, 27, 95, 97] or adjusting [88] loss functions.
4.1.3 Style-Driven Generation
The style-conditioned generation task aims to extract style information from given samples as conditions for controllable generation.
Similar to the approach of tuning-based subject-driven methods, StyleDrop [98] employs the fine-tuning of adapters [191, 201] on Muse [198] to tailor the model to specific style conditions and further proposes a prompt engineering technique to construct training data that effectively separates subject cues from style. Concurrently, several methods are exploring the incorporation of encoders designed to generate style-related embeddings for conditional generation [99, 100]. In addressing the issue of content borrowing from style references, ArtAdapter [100] introduces an innovative Auxiliary Content Adapter (ACA), which is designed to furnish the UNet with essential content cues, thereby ensuring that the model maintains a focus on style elements.
Besides, several approaches attempt to establish training-free frameworks for style-consistent image generation [101, 102]. For instance, StyleAligned [101] is designed to produce a series of images that adhere to a given reference style. This method introduces a novel attention sharing mechanism within the self-attention layers, which facilitates the interaction between the features of individual images and those of an additional reference image. Such a design enables the generation process to consider and incorporate style elements from multiple images simultaneously. Additionally, StyleAligned enhances the alignment of style attributes by normalizing both queries and keys using the Adaptive Instance Normalization (AdaIN) [202], further refining the style consistency across the generated images.
4.1.4 Interaction-Driven Generation
The interaction-conditioned generation task is specifically designed to learn and generate interaction-related concepts, such as human actions and human-object interactions (HOI). Essentially, this task centers around the novel idea of using a “verb” as the conditioning element.
For action-driven image generation, Huang et al. [189] proposes an Action-Disentangled Identifier (ADI) to decouple subject identity and action for improved action condition learning. To block the inversion of action-agnostic features, ADI extracts the gradient invariance from the constructed sample triples and masks the updates of irrelevant channels, which effectively ensures that the action condition is embedded in text embedding.
Moreover, Reversion [103] has been developed to comprehend relationships depicted in sample images, such as in scenarios where ’object A <is painted on>object B,’ with <is painted on>serving as the personalized condition. The method introduces a novel relation-steering contrastive learning mechanism, uniquely utilizing prepositions as positive samples to accurately guide the relational prompt, while other words are treated as negative samples. Additionally, Reversion employs a relation-focal importance sampling technique, which prioritizes the selection of samples with higher levels of noise during training, which facilitates the model’s learning of high-level semantic relationships.
Tian et al. [111] introduce the InteractDiffusion model to encapsulate human-object interaction (HOI) information for controllable generation. Central to their methodology is the construction of triplet labels encompassing a person, an action, and an object, along with corresponding bounding boxes, which are tokenized by interaction embeddings in InteractDiffusion to learn and represent the intricate relationships between these subjects.
4.1.5 Image-Driven Generation
The image-conditioned generation task aims to generate a similar image from multiple perspectives (e.g., content and style) by using an exemplar image as the prompt.
unCLIP [38] is an early work to explore using image prompt for image generation, proposing a two-stage model: a prior that generates a CLIP image embedding given a text caption, and a decoder that generates an image conditioned on the image embedding. Xu et al. [113] expand the existing single-flow diffusion pipeline into a multi-task multimodal network, dubbed Versatile Diffusion (VD), that handles multiple flows of text-to-image, image-to-text, and variations in one unified model. Xing et al. [114] introduce the Prompt-Free Diffusion to discard text with image in a text-to-image model, only using visual inputs to generate new images. They propose a Semantic Context Encoder (SeeCoder), consisting of a backbone encoder, a decoder, and a query transformer [195], to encode exemplar image. During inference, the SeeCoder will replace the CLIP text encoder in Stable Diffusion to take the reference image as input. IP-Adapter [116] decouples cross-attention mechanism that separates cross-attention layers for text features and image features to achieve image prompt capability for the pretrained text-to-image diffusion models. While these methods straightforwardly use images as prompts, ViscoNet [93] use segment person images to provide fashion reference for human-centric generation.
4.1.6 Distribution-Driven Generation
The distribution-conditioned generation task is designed to understand and learn from the data distribution of multiple exemplar images, with the aim of generating a variety of results reflective of this distribution. This approach is distinct from subject-centric personalization, as it focuses on adapting text-to-image models to generate broader, more abstract concepts or categories rather than individual subjects.
Cao et al. [119] introduce the Guidance-Decoupled Personalization framework, designed for generating specific concepts (e.g., faces) with a high degree of fidelity and editability. This framework uniquely decouples the conditional guidance into two distinct components: concept guidance and control guidance. The concept guidance component is specifically trained to steer the sampling process in a manner that adheres to the underlying data distribution, thereby ensuring the accurate generation of the reference concept. Moreover, DreamDistribution [120] is proposed for learning a prompt distribution, keeping a set of learnable text embeddings to model their distribution at CLIP text encoder feature space. Then a reparameterization trick is utilized to sample from this distribution and update the learnable embeddings.
4.2 Spatial Control
Since text is challenging to represent structure information, i.e. position and dense label, controlling text-to-image diffusion methods with spatial signals, e.g. layout[10, 203], human pose[204, 205, 206, 207], human parsing[208, 209, 210, 211, 212], and segmentation mask[213, 214], is a significant research field in diffusion models. Within this context, we commence with a brief overview of some unified methods for spatial control, followed by a more detailed exploration into various specific categories of structure, such as bounding boxes and keypoints.
4.2.1 Spatial-Conditional Score Prediction
In the domain of spatial-conditional score prediction, methods are developed to model $\hat{\epsilon}_{\theta}(x_{t},c_{text},c_{spatial},t)$ with the aim of generating results aligned with a given spatial condition $c_{spatial}$. We here overview model-based and training-free spatial-conditioned score prediction methods, since tuning-based methods [73] conceptualize structural conditions (e.g., layout) more abstractly and do not explicitly utilize spatial conditions.
• Model-based Score Prediction. ControlNet [190] stands out among generalized spatial controlling methods, gaining recognition as a seminal work and wining the prestigious Marr Award in 2023. Different significantly from methods that simply tune the parameters of the original diffusion model[87, 45], ControlNet introduces an innovative architecture by incorporating an additional encoder copy within the UNet structure. This added encoder is connected with the original UNet layers through the proposed “zero convolutions” to prevent overfitting and catastrophic forgetting. The simplicity and adaptability of ControlNet’s architecture have not only proven effective but also led to its widespread adoption as a baseline in numerous subsequent studies [186, 176, 136, 215, 216, 82, 217, 218]. Similarly, T2I-Adapter [219] is proposed to align internal knowledge in text-to-image diffusion models and external control signals. SCEdit [147] propose an efficient generative tuning framework, which integrates and edits skip connections using a lightweight tuning module named SC-Tuner.
While ControlNet [190] necessitates training distinct models for each type of controlling signal, some researchers have pursued the development of more generalized methods capable of handling a variety of spatial signals [215, 176, 115, 216]. To address this challenge, Qin et al. [215] introduce a task-aware HyperNet designed to modulate diffusion models for adaptability to different types of conditions, named UniControl. In this approach, conditions are encoded using a mixture of experts (MOE) adapter. Simultaneously, task instructions are transformed into task embeddings through the task-aware HyperNet, which are integrated to zero convolution for precise modulation of the condition features’ injection into the model.
In the arena of layout-conditioned score prediction, various innovative approaches have been introduced [123, 126, 128, 129, 122, 136, 141, 144, 111, 132, 143]. GLIGEN [123] employs grounded language as the basis for generation, embedding this grounding information into new trainable layers through a gated mechanism, thus enabling more controlled generation. In addition, SpaText [122] constructs spatio-textual representation by introducing CLIP image embeddings, where they stack these object embeddings in the same shapes and positions of the segments to control layout. Besides, some works focus on the face domain, synthesizing face images under face parsing condition [125, 84, 148, 148].
Several approaches have been developed to jointly denoise the spatial structure conditions for enhanced spatial control. JointNet [138], an extension of a pre-trained text-to-image diffusion model, introduces a new branch for dense modalities such as depth maps, where a duplicate of the original network is intricately connected with the RGB branch, facilitating complex interactions between different modalities. Additionally, Liu et al. [139] propose the Latent Structural Diffusion Model, which innovatively denoises depth and surface normal along with the RGB image synthesis.

### v2

@@ 3.1.1 Tuning-based Conditional Score Prediction
As the fundamental input of text-to-image diffusion models, text plays a central role in adapting these models to user-specific requirements. Textual Inversion (TI)[68] introduces an innovative idea by embedding user-provided concepts into new “words” in the text embedding space. This expands the tokenizer’s dictionary and optimizes additional tokens through a denoising process applied on provided images. DreamBooth[221] adopts a similar approach but encodes concepts with low-frequency words (i.e., sks), and additionally updates the UNet parameters with a class-specific prior preservation loss to enhance output diversity. Together, the simplicity and adaptability of TI and DreamBooth have established them as foundational frameworks for many subsequent tuning-based methods. Building on these foundations, Custom Diffusion [130] examines weight deviations during fine-tuning and identifies cross-attention layer parameters—particularly the key and value projections ($W^{k}$ and $W^{v}$)—as pivotal. It thus narrows updates to these projections and augments the process with extra text tokens and a regularization loss.
To further improve textual inversion, recent works explore layer-specific distinctions in UNet [262, 328, 114], refine embedding initialization strategies [189], and sampling distribution [332]. These methods apply distinct text embeddings across layers to capture finer variations. In contrast, CatVersion [335] moves away from tuning text embeddings or UNet parameters, and instead learns concatenated embeddings within the dense feature space of the text encoder. This design proves effective for capturing subtle differences between a personalized concept and its base class, thereby helping preserve prior knowledge.
• Parameter-Efficient Fine-Tuning (PEFT). Beyond full fine-tuning, parameter-efficient methods (PEFT)[96, 98, 258, 26] have become increasingly important in personalization [294]. Among them, Low-rank Adaptation (LoRA)[98] has been widely adopted in various personalization pipelines [221, 31, 75, 236, 78, 175]. Xiang et al. propose ANOVA [289], which employs adapters [96] and demonstrates that placing them after the cross-attention block notably boosts performance. COMCAT [294] develops an efficient ViT [60] compression method based on model factorization. Similarly, DiffuseKronA [175] introduces a Kronecker product-based adaptation module that surpasses LoRA-DreamBooth in parameter efficiency and stability, offering consistent high-quality generation with greater interpretability. To support research and evaluation in this area, LyCORIS [308] provides a comprehensive open-source library [^1] https://github.com/KohakuBlueleaf/LyCORIS covering numerous PEFT methods (e.g., LoRA, LoHa, DyLoRA [258]) and a framework for their systematic assessment, thus promoting the progress of diffusion model personalization.
• Condition Disentanglement. A further challenge in introducing novel conditions lies in disentangling the desired concept from confounding inputs. Many studies [28, 329, 78, 237, 6, 31, 19, 141, 178, 116, 326, 347, 324, 109, 224, 110] observe that irrelevant information—such as background context or co-occurring objects—tends to be entangled with the target concept. To address this, several works [6, 115, 223] employ explicit masks to isolate object regions. In this direction, Disenbooth [31] and DETEX [19] reduce the impact of backgrounds, with DETEX further decoupling pose from subject appearance. PACGen [141] instead applies aggressive data augmentation, altering the size and position of concepts to help separate spatial cues from the core identity.
Other approaches disentangle conditions by adjusting fine-tuning strategies. DreamVideo [282] separates subject learning and motion learning, while B-LoRA [66] jointly optimizes LoRA weights of two blocks to implicitly decouple style and content. ReVersion [107] introduces a relation-steering contrastive learning scheme to capture object relationships more effectively.
Training techniques can also be exploited to reduce condition entanglement [122, 104, 288]. Selective Information Description [122] leverages a VLM to produce refined text descriptions, ensuring the model emphasizes target objects over contextual biases. Similarly, U-VAP [288] adopts a decoupled self-augmentation strategy, where an LLM generates paired target and non-target prompts, facilitating dual concept learning to decouple conditions.
• Prior Preservation. Another important challenge in fine-tuning diffusion models is the preservation of prior knowledge. Without careful design, models risk overfitting to narrow concepts, sacrificing generality and controllability. To address this, prior preservation techniques have been widely explored [221, 130, 253, 207, 83, 279, 314, 204]. Perfusion [253] mitigates overfitting by locking cross-attention keys to prior categories and applying a gated rank-1 concept update. SVDiff [83] regulates singular values in weight matrices to reduce risks such as language drift, while OFT [207] introduces orthogonal fine-tuning to preserve semantic capacity by maintaining hyperspherical energy. Together, these methods help balance fidelity to new data with the retention of broad generative ability.
• Training Techniques. Beyond the strategies above, alternative training techniques have been explored to improve efficiency, reduce computational overhead, and enhance performance [260, 63, 3, 300, 88, 335, 325]. For example, DVAR [260] proposes a variance-based early stopping criterion to replace unreliable convergence metrics, thereby accelerating training. Gradient-Free Textual Inversion [63] divides optimization into dimensionality reduction and non-convex gradient-free search, achieving faster convergence with minimal performance loss. MATTE [3] investigates the roles of timesteps and UNet layers for different concept categories, aiming for broader adaptability.
Another notable direction is the introduction of auxiliary losses to boost generation performance [281, 112, 276, 287]. TokenCompose [281] improves multi-object composition and photorealism through token-wise consistency losses between images and segmentation maps. Similarly, CoMat [112] addresses text-image misalignment via a concept activation loss and enhances visual quality with adversarial loss.
Lastly, data-centric strategies have been proposed to enhance the training process. COTI [300] improves Textual Inversion through active and controllable data selection, while He et al. [88] generate text- and image-level regularization datasets to better preserve model generalization.
• Inference Techniques. In addition to training methods, several approaches enhance controllability during inference by modulating cross-attention to better align outputs with target conditions [342, 81, 182, 346]. For example, MagicTailor [342] combats semantic pollution and imbalance with dynamic masked degradation and dual-stream balancing. OMG [127] provides an occlusion-friendly framework that adopts a two-stage sampling process—first generating layout and visual comprehension for occlusion handling, then applying noise blending to integrate concepts—leading to superior identity preservation and visual harmony.
Additionally, DreamBlend [211] addresses the trade-off between prompt fidelity, subject fidelity, and diversity by leveraging multiple checkpoints and combining their strengths through cross-attention guidance. Prompt-aligned personalization [5] improves complex text alignment via score distillation sampling while supporting multi-/single-shot scenarios, subject composition, and reference-guided generation.
Figure 4: Illustration of adapter-based conditional score prediction.
3.1.2 Adapter-based Conditional Score Prediction
To eliminate test-time tuning cost, a class of methods introduces an additional encoder $E$ that maps novel conditions to feature embeddings and feeds them into the noise predictor. The conditional score prediction then reads
$\tilde{\epsilon}(x_{t},c_{text},c_{novel},t)$ $=\epsilon_{\theta^{*}}\!\big(x_{t},\,c_{text},\,E(c_{novel}),\,t\big),$ (17)
where $\theta^{*}$ and $E$ are learned offline and kept fixed at inference. Fig. 4 gives a schematic overview.
Adapter-based variants primarily differ by the condition families they target (e.g., geometry, style, layout, controls), which dictate how the introduced adapter $E(\cdot)$ interfaces with the backbone (e.g., concatenation to token embeddings, cross-attention keys/values, or feature-wise affine modulation). In what follows, we group methods by task type and provide a systematic review; task definitions are summarized in Sec. 2.4.
• Spatial Control. ControlNet [323] stands out among generalized spatial controllers, earning recognition as a seminal work and winning the Marr Award in 2023. Distinct from approaches that simply fine-tune the base diffusion model parameters [222, 130], ControlNet introduces an auxiliary encoder branch mirroring the U-Net and couples it to the original layers via zero convolutions to mitigate overfitting and catastrophic forgetting. Owing to its simplicity and adaptability, ControlNet has proved effective and has become a widely adopted baseline in subsequent studies [99, 111, 206, 302, 313, 293, 171, 272]. Similarly, T2I-Adapter [179] aligns internal knowledge in text-to-image diffusion models with external spatial control signals.
While ControlNet [323] requires training a separate model for each control type, subsequent work seeks general controllers that handle diverse spatial signals [206, 99, 336, 302, 65, 148, 341]. Qin et al. [206] propose UniControl, a task-aware HyperNetwork that modulates the diffusion backbone across condition types: conditions are encoded via a mixture-of-experts (MoE) adapter, while task instructions are embedded by the HyperNet and injected through zero-conv gating to precisely regulate how condition features enter the model. Meta ControlNet [302] adapts the pretrained ControlNet to another condition domain by meta learning, apparently reducing the learning steps. In parallel, Ctrl-Adapter [148] and X-Adapter [215] transfer diverse controls to arbitrary diffusion backbones by adapting pretrained ControlNets.
For layout-conditioned generation, many methods arrange regions and bind them to textual concepts [142, 82, 299, 340, 7, 111, 273, 203, 94, 36, 263]. GLIGEN [142] grounds language with structured inputs and injects the grounding via gated trainable layers, enabling controllable placement. SpaText [7] builds a spatio-textual representation by stacking CLIP-derived object embeddings in masks at their target positions to enforce layout. Related efforts focus on the face domain, synthesizing faces under face-parsing constraints [46, 73].
From the objective perspective, several approaches jointly denoise dense spatial conditions to improve alignment. JointNet [319] augments a pretrained T2I backbone with a dense-modality branch (e.g., depth) that is tightly coupled to the RGB branch, enabling rich cross-modality interactions. Liu et al. [159] propose a Latent Structural Diffusion Model that co-denoises depth and surface normals alongside RGB synthesis. Complementarily, adding auxiliary spatial-consistency losses further enhances control [138, 126, 144]. Specifically, ControlNet++ [138] enforces pixel-level cycle consistency between outputs and controls, while Li et al. [144] introduce a segmentation-based discriminator to provide explicit spatial feedback.